Find the finite population correction factor given the following: 1. N=200, n=20 2. N=250, n=10 3. N=2000, n=15 4. N=2500, n=20 5. N=200, n=20

1
Expert's answer
2022-02-24T17:54:55-0500

Finite population correction factor 


"FPC=\sqrt{\dfrac{N-n}{N-1}}"

1.


"FPC=\sqrt{\dfrac{200-20}{200-1}}=0.951064"

2.


"FPC=\sqrt{\dfrac{250-10}{250-1}}=0.981761"


3.


"FPC=\sqrt{\dfrac{2000-15}{2000-1}}=0.996492"



4.


"FPC=\sqrt{\dfrac{2500-20}{2500-1}}=0.996191"

5.


"FPC=\sqrt{\dfrac{200-10}{200-1}}=0.977125"